OpenStudy (anonymous):

ten times the square of a nonzero number is equal to forty times the number. what is the number?

OpenStudy (anonymous):

\[10*x^2=40x\] the question's statement \[x^2-4x=0\] simplify \[x(x-4)=0\] factor out an x x=0,x=4 the two solutions

OpenStudy (anonymous):

wait so whats the answer? like what do I put in? it is not multiple choice either @GaryCollins

OpenStudy (anonymous):

State that x can be 0 or 4

OpenStudy (anonymous):

But the answer would be that it is 0 AND 4